OpenStudy (anonymous):

Complex numbers are written in the form a + bi. What does i represent ? would that be radical -1 ?

OpenStudy (anonymous):

Yes.. i =\[i=\sqrt{-1}\] because when you square both sides, you get \[i ^{2}=\left( \sqrt{-1} \right)^{2}=-1\]

OpenStudy (anonymous):

So when you look at the expression \[\sqrt{-25}=\sqrt{-1}\sqrt{25}=i \times5=5i\]